The SAT math exam has mean of 451 and standard deviation 17. Jimmy scored at the D9 level. What was his raw score?
Answer by stanbon(75887)

The SAT math exam has mean of 451 and standard deviation 17. Jimmy scored at the D9 level. What was his raw score?
---
Find the z-value with a 90% left tail.
invNorm(0.9) = 1.2816
=======================
Find the corresponding score value
x = z*s +u
x = 1.2816*17+451 = 472.79
=====================================
Cheers,
Stan H.
========================